OpenStudy (ellie202000):

Please help!!! I will medal the best response!!! Given the system of equations, what is the value of the y-determinant? 3x+y-10=0 4x-y-4=0 A)-28 B)-14 C)28 What did I do wrong and how do I fix it? Here is how I did it...

OpenStudy (ellie202000):

The answers A, but how? I did the problem twice and got \[\frac{ 52 }{ 7 }\] Both times. Here's how I did it: I first convered them to 4x+(-y)=4 3x+y=10 Then:  Then:  So after all of that it ends up to be \[7 \frac{ 3 }{ 7}\] What did I do wrong?